OpenStudy (oraclethinktank):

Help please. I need to graph the Quadratic Equation y = -x^2 - 4x - 3

OpenStudy (oraclethinktank):

So far I have a= -x^2 b= -4x c= -3 Following the rule of x= -b/2a x= 4x/2(-x^2) = 4x/-2x^2 = - 2/x

OpenStudy (oraclethinktank):

Is the line of symmetry - 2/x?

OpenStudy (calculusxy):

Remember that a, b, and c are actually the numbers that are the coefficients. For example in 2x^2 + 4x + 6 = 0 a = 2 b = 4 c = 6

OpenStudy (oraclethinktank):

So for my equation, they would be a=2 b=-4 c=-3

OpenStudy (calculusxy):

In -x^2 - 4x - 3 a = -1 b = -4 c = -3

OpenStudy (oraclethinktank):

So with that, using x=-b/2a would be 4/2?

OpenStudy (oraclethinktank):

Or 4/-2, rather.

OpenStudy (oraclethinktank):

Which in turn means that the Line of Symmetry for this equation is -2 since 4/-2 = -2

OpenStudy (calculusxy):

The axis of symmetry is x = -(b)/2a in this case you would have x = -(-4)/2(-1) x = 4/-2 x = -2

OpenStudy (oraclethinktank):

So I have that correct, What would come next for graphing the equation?

OpenStudy (calculusxy):

Find the vertex by plugging back -2 into the quadratic equation to get the y value.

OpenStudy (calculusxy):

Then make a table with